# Cost-aware Stopping for Bayesian Optimization

### v1

@@ 3.1 A Theoretical Guarantee on Cost up to Stopping
Let $X$ be compact, and let $f:X\to\mathbb{R}$ be a random function with constant mean $\mu$, meaning $\mathbb{E}[f(x)]=\mu$ for all $x\in X$. Consider a Bayesian optimization algorithm which begins at some initial point $x_{1}\in X$ with cost $C=c(x_{1})$, acquires additional points using either PBGI or LogEIPC, and uses the PBGI/LogEIPC stopping rule. Let $\tau=\min_{t\geq 1}\{\sup_{x\in X}\alpha_{t}^{\mathrm{LogEIPC}}(x)\leq 0\}$ be its stopping time, and let $U=\mu-\mathbb{E}[\min_{x\in X}f(x)]<\infty$. Then we have
$\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq C+U.$ (12)
Here, $U$ serves as an upper bound on the expected cumulative improvement $\mathbb{E}[y_{1:1}^{*}-y_{1:\tau}^{*}]$. When $f\sim\mathcal{GP}(\mu,K)$, the term $U$ on the right-hand side can be further upper bounded using classic results on the expected supremum of a Gaussian process; see, for example, Lifshits [14, Theorem 10.1].
A proof of Theorem 1 is given in Appendix A. The proof also indicates that the theorem extends to acquisition functions beyond PBGI and LogEIPC that ensure sufficient expected improvement relative to cost before the stopping time $\tau$; see condition (17). This result demonstrates that the PBGI stopping rule avoids excessive spending not just in Pandora’s Box [25], but also Bayesian optimization with general correlations. As discussed below, Theorem 1 has another important implication for settings where there is no direct conversion between objective function values and evaluation costs.
3.2 Practical Implementation Considerations for the PBGI/LogEIPC Stopping Rule
Choosing a cost scaling factor $\lambda$.
In practice, evaluation costs often come in units that differ from how benefit is measured—for example, run time versus monetary costs. To reconcile this, we multiply costs by a scaling factor $\lambda>0$—that is, we use cost function $\lambda c(x)$ instead of $c(x)$, in order to ensure that costs $\lambda c(x)$ and objective values $f(x)$ have the same units. For example, if one is willing to spend $1000$ seconds of time to improve the objective by $0.01$, then one should set $\lambda=10^{-5}$.
If one knows the unit conversion explicitly—for instance, the hourly rate of a compute cluster and the business value of improved model performance—then the scaling factor $\lambda$ can be set directly. Otherwise, Theorem 1 gives a principled heuristic for choosing $\lambda$. When applied in the scaled-cost setting, Theorem 1 bounds the expected cumulative unscaled cost by $C+U/\lambda$. This implies that, given an expected cost budget $B$, one can safely use $\lambda=U/(B-C)$; under an almost-sure budget $B$, using PBGI-D [25], one can initialize $\lambda_{0}=U/(B-C)$, ensuring that the adaptive decay is triggered quickly in expectation. This provides a way to automatically set the hyperparameters $\lambda$ and $\lambda_{0}$ when using PBGI and PBGI-D, respectively—removing the need for manual tuning.
@@ 3.2 Practical Implementation Considerations for the PBGI/LogEIPC Stopping Rule
In practice, evaluation costs are often not known in advance. They can be modeled in either: (i) deterministically, using domain knowledge, for instance, proportional to the number of model parameters in hyperparameter tuning, or (ii) stochastically, via Gaussian process over log costs. For the latter, Astudillo et al. [2] proposed a general framework for handling unknown costs, which can also be applied to stopping rules. For the LogEIPC stopping rule, following their Proposition 2, the $c(x)$ in (10) can be replaced with
$\mathbb{E}[1/c(x)]^{-1}=\exp\left\{\mu_{\ln c}(x)-(\sigma_{\ln c}(x))^{2}/2\right\}.$ (13)
Alternatively, one could also replace it with
$\mathbb{E}[c(x)]=\exp\left\{\mu_{\ln c}(x)+(\sigma_{\ln c}(x))^{2}/2\right\}.$ (14)
The difference in sign before the variance term reflects how each formulation responds to predictive uncertainty: (13) encourages exploration more than (14). For PBGI, under the unknown-cost setting, $c(x)$ in (5) should be replaced by $\mathbb{E}[c(x)]$ using (14), reflecting the manner in which costs enter the respective root-finding problem. This means that in unknown-cost settings, the equivalence between the PBGI and LogEIPC stopping rules holds when replacing $c(x)$ by (14), but not by (13).
@@
Although our stopping rule has theoretical guarantees, in practice, it—like other adaptive stopping rules—can still suffer from spurious stops caused by two main sources which we now discuss. First, early in the optimization process, the Gaussian process’ model parameters often fluctuate significantly as new data points are collected, causing unstable stopping signals. To mitigate this, we enforce a stabilization period consisting of the first several evaluations, during which we do not allow any stopping rule to trigger. Second, imperfect optimization of the acquisition function—which is especially common in higher-dimensional search spaces—can lead to misleading stopping signals. To handle this, we use a debounce strategy, requiring the stopping rule to consistently indicate stopping over several consecutive iterations before stopping optimization.
@@ 4 Experiments
To evaluate our proposed PBGI/LogEIPC stopping rule, we design three complementary sets of experiments that progressively test its performance. First, we consider an idealized, low-dimensional Bayesian regret setting in which the Gaussian process model exactly matches the true objective. This controlled environment allows us to isolate the effect of different stopping rules without interference from modeling or optimization errors. Then, we move to an 8-dimensional Bayesian regret setting where each acquisition-function minimization must be approximated via a numerical optimizer. Finally, we test in practical scenarios with potential model mismatch, using the LCBench hyperparameter tuning benchmark and the NATS neural architecture size search benchmark. In each case, we compare combinations of acquisition functions and stopping rules, which we describe next.
@@
We consider four common acquisition functions that were discussed in Section 2: log expected improvement per cost (LogEIPC), Pandora’s Box Gittins Index (PBGI), lower confidence bound (LCB), and Thompson sampling (TS)—chosen for their competitive performance, computational efficiency, and close connections to existing stopping rules.
@@
We compare the proposed PBGI/LogEIPC stopping rule against several stopping rules from prior work: UCB-LCB [17], LogEIPC-med [11], SRGap-med [11], and PRB [24]. We also include two simple heuristics used in practice: Convergence and GSS. UCB-LCB stops once the gap between upper and lower confidence bounds falls below a configurable threshold $\theta$. LogEIPC-med stops when the log expected improvement per cost drops beneath $\log(\eta)$ plus the median of its initial $I$ values. SRGap-med stops when the simple-regret gap falls below $\chi$ times the median of its initial $I$ values. PRB triggers once a probabilistic regret bound satisfies the regret tolerance $\varepsilon$ and confidence $\delta$ parameters. Convergence stops as soon as the best observed value remains unchanged for $w$ iterations. GSS stops if the recent improvement is less than $\phi\times\mathrm{IQR}$ over the past $w$ trials where $\mathrm{IQR}$ denotes the inter-quartile range of current observations. Finally, we include Hindsight, which, for each trial, selects the stopping time that yields the lowest cost-adjusted simple regret in hindsight, thus providing a lower bound on achievable performance.
In our experiments, unless specified otherwise, we follow parameter values recommended in the literature, and set $\theta=0.01$, $\eta=0.01$, $\chi=0.01$, $I=20$, $\epsilon=0.05$, $\delta=0.05$, $w=5$, and $\phi=0.01$. Each experiment is repeated with 50 random seeds to assess variability, and we report the mean with error bars (2 times the standard error) for each stopping rule. Each trial, in the sense of a run with a distinct random seed, is capped at a fixed number of iterations; if a stopping rule is not triggered within this limit, the stopping time is set to the cap. Details are provided in Appendix B.
@@ 4.1 Bayesian Regret
In the one-dimensional experiments, we perform an exhaustive grid search over $[0,1]$ to isolate the effect of stopping behavior from numerical optimization. In the 8-dimensional setting, due to instability from higher dimensionality, we apply moving average with a window size of 20 when computing the PBGI/LogEIPC stopping condition. See Figure 6 in Appendix B for an illustration. We omit PRB in the 8D experiment due to its high computational overhead. More details on our experiment setup, and computational considerations are provided in Appendix B and Appendix C.
Figure 2: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in $1$-dimensional Bayesian regret experiments. Objective functions are sampled from a Gaussian process with a Matérn-5/2 kernel, and a linear cost function is applied with scaling factors $\lambda=0.1,0.01,0.001$. The matched PBGI/LogEIPC combinations perform best at $\lambda=0.1$ and $0.01$, and although slightly suboptimal at $\lambda=0.001$ compared to PRB paired with the PBGI or LogEIPC acquisition function, still achieves cost-adjusted regret close to the hindsight optimal.
Figure 3: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in $8$-dimensional Bayesian regret experiments. Objective functions are sampled from a Gaussian process with a Matérn-5/2 kernel, using three different cost functions scaled by $\lambda=0.01$. The PBGI/LogEIPC stopping rule consistently yields the lowest cost-adjusted regret that is close to hindsight optimal, when paired with LogEIPC, PBGI, and LCB.
@@
Figure 2 shows a comparison of cost-adjusted regret for acquisition function and stopping rule pairings with different cost-scaling factors $\lambda$ in the 1-dimensional setting. From the plot, our PBGI/LogEIPC stopping rule pairing with LogEIPC or PBGI acquisition function delivers competitive performance for each $\lambda$, and is particularly strong in handling high cost scenarios—here, those with large $\lambda$.
@@
In the 1-dimensional setting, the Bayesian optimization problem is relatively straightforward and all acquisition functions attain nearly identical hindsight optima, independent of cost scale or cost type. In the 8-dimensional experiments, however, clear gaps emerge: Figure 3 compares cost-adjusted regret under uniform, linear, and periodic cost regimes. Under uniform and linear costs, LogEIPC, PBGI, and LCB perform similarly and substantially outperform Thompson Sampling (TS), while in the periodic case, LogEIPC and PBGI outperform LCB and TS. In every case, combining our stopping rule with the LogEIPC, PBGI, or LCB acquisition function yields cost-adjusted regret nearly matching the hindsight optimal. This shows our PBGI/LogEIPC stopping rule to be relatively robust against challenges in acquisition function optimization. Further discussions of our experiments across all cost types and cost-scaling factors can be found in Appendix C.
4.2 Empirical Benchmarks: Hyperparameter Optimization and Neural Architecture Search
@@
We now consider two empirical benchmarks based on use cases from hyperparameter optimization and neural architecture size search: LCBench [27] and NATS-Bench [7].
@@
LCBench provides training data of 2,000 hyperparameter configurations evaluated on multiple OpenML datasets. In the unknown-cost experiments, we consider the full training (at 200 epochs) time as the cost. For the known-cost setting, we observe that training time to scale approximately linearly with the number of model parameters. Based on a linear regression fit (see Appendix B), we adopt $10^{-3}\times p$, where $p$ is the number of model parameters, as a proxy of training time.
@@
NATS-Bench defines a search space of 32,768 neural architectures varying in channel sizes across layers, evaluated on cifar10-valid, cifar100, and ImageNet16-120. As in LCBench, for the known-cost experiments, we approximate the full training + validation time at 90 epochs using $\alpha F+\beta$, where $F$ is the number of floating point operations; see Appendix B for details.
@@
We report the mean and error bars (2 times standard error) of the cost-adjusted simple regret, where we consider the evaluation costs to be some cost-scaling factor $\lambda$ (see Section 3.2) multiplied with cumulative runtime. Simple regret is computed as the difference between (a) test error of the configuration with best validation error at the stopping time and (b) the best test error over all configurations. We present the results using proxy runtime here, deferring the results using actual runtime—an unknown-cost scenario—to Appendix C. Additionally, we report the means and error bars of cumulative runtime and test error pairs at stopping in Appendix C.
@@
Figure 4 and Figure 5 present cost-adjusted regret for combinations of acquisition function and stopping rule on classification tasks from LCBench and NATS, with cost-scaling factor $\lambda=10^{-4}$. Additional results under different cost-scaling factors and the unknown-cost setting are provided in Appendix C, along with visual comparisons between adaptive and fixed-iteration stopping. Overall, our PBGI/LogEIPC stopping rule consistently performs strongly in terms of cost-adjusted simple regret, particularly when paired with the PBGI acquisition function. We also report, in Table 1 of Appendix C, how often each stopping rule fails to trigger within the 200-iteration cap: baselines such as SRGap-med and UCB-LCB often fail to stop early—particularly on NATS—whereas our PBGI/LogEIPC stopping rule reliably stops before the cap.
Figure 4: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on LCBench. The objective is to minimize validation error on three classification tasks, with proxy runtime used as the evaluation cost, scaled by $\lambda=10^{-4}$. The matched PBGI combination consistently performs best on higgs and volkert, and second-best on Fashion-MNIST, closely approaching the hindsight optimal across all tasks. Notably, our stopping rule also performs competitively when paired with LogEIPC or Thompson sampling.
@@
As we transition from model match to model mismatch, where the true objective function does not align perfectly with the Gaussian process prior, we find that the PBGI/LogEIPC stopping rule, when paired with the PBGI acquisition function, continues to deliver performance closest to the hindsight optimal. However, pairing the same stopping rule with the LogEIPC acquisition function becomes noticeably less competitive on LCBench datasets. This degradation appears to stem from the relative advantage of PBGI over LogEIPC in higher dimensions or misspecified settings: we illustrate this in more detail in Appendix C, which shows that the PBGI acquisition function maintains stronger performance under mis-specification, thus contributing to the improved overall cost-adjusted regret. These results highlight that, while our stopping rule is broadly robust, the choice of acquisition function remains an important consideration when facing model mismatch.
Figure 5: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on NATS-Bench. The objective is to minimize validation error on three image classification tasks, with proxy runtime used as the evaluation cost, scaled by $\lambda=10^{-4}$. Our stopping rule consistently approaches the hindsight optimal, though it is not always the top performer—GSS and Convergence can achieve lower cost-adjusted regret. In contrast, the remaining stopping rules are rarely triggered within the 200-iteration budget. See Table 1 in Section C.3 for details.

### v2

@@ 3.1 Theoretical Guarantees on Cost-Adjusted Simple Regret
Consider the setting, acquisition function, and stopping rule specified in Lemma˜1, and let $U=\mu-\mathbb{E}[\min_{x\in X}f(x)]<\infty$. Then the expected cost-adjusted simple regret is bounded by
$\mathbb{E}\left[y_{1:\tau}^{*}-\min_{x\in X}f(x)+\sum_{t=1}^{\tau}c(x_{t})\right]\leq\mu+C-\mathbb{E}\left[\min_{x\in X}f(x)\right]=\mathbb{E}\Big[\underbrace{y_{1}-\min_{x\in X}f(x)+c(x_{1})}_{\textnormal{cost-adjusted regret of "Immediate"}}\Big].$ (11)
3.1.1 Implication in the Budget-constrained Setting
We first note that in the discrete Pandora’s Box setting, under an expected budget constraint $B$, minimizing $\alpha_{t}^{\mathrm{PBGI}}(x)$ is Bayesian-optimal: it achieves the lowest expected simple regret, among all algorithms satisfying $\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq B$, and the cost function used in defining $\alpha_{t}^{\mathrm{PBGI}}(x)$ is $\lambda c(x)$ for some cost-scaling factor $\lambda$ which depends on $B$ (Xie et al., 2024, Theorem 2).
This result, at first, appear to be completely Pandora’s-Box-theoretic: it requires $X$ to be discrete and $f$ to be independent. In the more general correlated setting of cost-aware Bayesian optimization, however, Bayesian optimality of PBGI may no longer hold, and the relationship between $B$ and the choice of $\lambda$ is not immediately clear. Theorem˜2 helps bridge this gap: it provides an upper bound on the expected cumulative cost up to the stopping time, which in turn yields the following principled choice of $\lambda$ that ensures compliance with the budget constraint.
Corollary 3.
Consider the setting, acquisition function, stopping rule, and notation specified in Lemmas˜1 and 2, but with costs rescaled by a factor $\lambda>0$: both acquisition values and stopping conditions are computed using $\lambda c(\cdot)$. If the cost-scaling factor is set to $\lambda=\frac{U}{B-C}$, then the algorithm’s expected cumulative unscaled cost satisfies $\mathbb{E}\!\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq B$.
If this choice of $\lambda$ proves overly conservative—leading to underspending within the budget—it can be paired with the PBGI-D variant of Xie et al. (2024), which starts with $\lambda=\lambda_{0}$ and halves it each time stopping is triggered. Choosing $\lambda_{0}=U/(B-C)$ aligns the initial fixed-$\lambda$ phase with the budget in expectation, while ensuring that the adaptive decay is activated as designed. In Figure˜11 in Appendix˜D, we show PBGI-D with this choice of $\lambda_{0}$ is competitive.
3.2 Practical Implementation Considerations for PBGI/LogEIPC Stopping Rule
Expressing objective and cost in common units.
Evaluation costs are often measured in different units than the objective, such as time versus accuracy. To compare them directly, we rescale costs by a constant $\lambda>0$, the conversion rate between cost and objective improvement. For instance, if one is willing to spend $1000$ seconds to reduce test error by $0.01$, then $\lambda=10^{-5}$. Importantly, in the cost-per-sample setting we mainly study, $\lambda$ is set by the problem provider rather than tuned by the user, whereas the budgeted setting without a natural unit conversion is discussed in Section˜3.1.1.
@@ 3.2 Practical Implementation Considerations for PBGI/LogEIPC Stopping Rule
In practice, evaluation costs are often not known in advance. They can be modeled either (i) deterministically, using domain knowledge (e.g., proportional to model size in hyperparameter tuning), or (ii) stochastically, via a Gaussian process over log costs. In Section˜4, we present empirical benchmark results under both modeling approaches. For the stochastic approach, we follow Astudillo et al. (2021, Proposition 2) to model $\ln c(x)$ as a Gaussian process with posterior mean $\mu_{\ln c}$ and variance $\sigma^{2}_{\ln c}$ . To compute LogEIPC or PBGI, as well as their related stopping rules (including prior acquisition-based ones and ours), we replace $c(x)$ in Equations˜4, 5 and 9 with $\mathbb{E}[c(x)]=\exp(\mu_{\ln c}(x)+(\sigma_{\ln c}(x))^{2}/2)$. See Section˜D.3 for further discussion.
@@
Although our stopping rule has theoretical guarantees, in practice, it—like other adaptive stopping rules—can still suffer from spurious stops caused by two main sources which we now discuss. First, early in the optimization process, the Gaussian process model parameters often fluctuate significantly as new data points are collected, causing unstable stopping signals. To mitigate this, we enforce a stabilization period consisting of the first several evaluations, during which we do not allow any stopping rule to trigger. Second, imperfect optimization of the acquisition function—which is especially common in higher-dimensional search spaces—can lead to misleading stopping signals. To handle this, we use a debounce strategy, requiring the stopping rule to consistently indicate stopping over several consecutive iterations before stopping optimization.
@@ 4 Experiments
To evaluate our proposed PBGI/LogEIPC stopping rule, we design three complementary sets of experiments that progressively test its performance. First, we consider an idealized, low-dimensional Bayesian regret setting in which the Gaussian process model exactly matches the true objective. This controlled environment allows us to isolate the effect of different stopping rules without interference from modeling or optimization errors. Then, we move to a higher-dimensional Bayesian regret setting where each acquisition-function minimization must be approximated via a numerical optimizer. Finally, we test in practical scenarios with potential objective model mismatch, using the LCBench hyperparameter tuning benchmark and the NATS neural architecture size search benchmark. In each case, we compare combinations of acquisition functions and stopping rules, which we describe next.
@@
We consider four common acquisition functions that were discussed in Section˜2: log expected improvement per cost (LogEIPC), Pandora’s Box Gittins Index (PBGI), lower confidence bound (LCB), and Thompson sampling (TS)—chosen for their competitive performance, computational efficiency, and close connections to existing stopping rules.
@@
We compare the proposed PBGI/LogEIPC stopping rule against several stopping rules from prior work: UCB-LCB Makarova et al. (2022), LogEIPC-med Ishibashi et al. (2023), SRGap-med Ishibashi et al. (2023), and PRB Wilson (2024). We also include two simple heuristics used in practice: Convergence and GSS. UCB-LCB stops once the gap between upper and lower confidence bounds falls below a configurable threshold $\theta$. LogEIPC-med stops when the log expected improvement per cost drops beneath $\log(\eta)$ plus the median of its initial $I$ values. SRGap-med stops when the gap of the expected minimum simple regret falls below $\chi$ times the median of its initial $I$ values. PRB triggers once a probabilistic regret bound satisfies the regret tolerance $\epsilon$ and confidence $\delta$ parameters. Convergence stops as soon as the best observed value remains unchanged for $w$ iterations. GSS stops if the recent improvement is less than $\phi\times\mathrm{IQR}$ over the past $w$ trials where $\mathrm{IQR}$ denotes the inter-quartile range of current observations. Finally, we include two reference baselines: Immediate, which stops right after the initial evaluation (see Section˜3.1), and Hindsight, which, for each trial, selects the stopping time that yields the lowest cost-adjusted simple regret in hindsight, thus providing a lower bound on achievable performance.
In our experiments, unless specified otherwise, we follow parameter values recommended in the literature, and set $\theta=0.01$, $\eta=0.01$, $\chi=0.01$, $I=20$, $\epsilon=0.1$ for Bayesian regret experiments, $\epsilon=0.5\%$ of the best test error for empirical experiments, $\delta=0.05$, $w=5$, and $\phi=0.01$. Each experiment is repeated with 50 random seeds to assess variability, and we report the mean with error bars (2 times the standard error) for each stopping rule. Each trial, in the sense of a run with a distinct random seed, is capped at a fixed number of iterations; if a stopping rule is not triggered within this limit, the stopping time is set to the cap. Details are provided in Appendix˜C.
@@ 4.1 Bayesian Regret
In the 1-dimensional experiments, we perform an exhaustive grid search over $[0,1]$ to isolate the effect of stopping behavior from numerical optimization. In the 8-dimensional setting, due to instability from higher dimensionality, we apply moving average with a window size of 20 when computing the PBGI/LogEIPC stopping condition. See Figure˜5 in Appendix˜C for an illustration. More details on our experiment setup, and computational considerations are provided in Appendices˜C and D.
@@
Figure˜2 shows a comparison of cost-adjusted regret for acquisition function and stopping rule pairings under different cost-scaling factors $\lambda$ in the 1-dimensional setting. From the plot, our PBGI/LogEIPC stopping rule pairing with LogEIPC or PBGI acquisition function delivers competitive performance for each $\lambda$, and is particularly strong in handling high-cost scenarios—those with large $\lambda$.
@@
In the 1-dimensional setting, the Bayesian optimization problem is relatively straightforward and all acquisition functions attain nearly identical hindsight optima, independent of cost scale or cost type. In the 8-dimensional experiments, however, clear gaps emerge: Figure˜3 compares cost-adjusted regret under uniform, linear, and periodic cost regimes. Under uniform and linear costs, LogEIPC, PBGI, and LCB perform similarly and substantially outperform TS, while in the periodic case, LogEIPC and PBGI outperform LCB and TS. In every case, combining our stopping rule with the LogEIPC, PBGI, or LCB acquisition function yields cost-adjusted regret nearly matching the hindsight optimal. This shows our PBGI/LogEIPC stopping rule to be relatively robust against challenges in acquisition function optimization. Further discussions of our experiments across all cost types and cost-scaling factors can be found in Appendix˜D.
Figure 2: Cost-adjusted simple regret across acquisition–stopping rule pairs in 1D Bayesian regret setting. Objectives are sampled from a GP with Matérn-5/2 kernel, with a linear cost function scaled by $\lambda=0.1,0.01,0.001$. Matched PBGI/LogEIPC pairs perform best at $\lambda=0.1,0.01$, and remain near-hindsight-optimal at $\lambda=0.001$, though slightly outperformed by PRB + PBGI/LogEIPC acquisition. The Immediate baseline is omitted at $\lambda=0.001$ due to its significantly higher regret compared to other stopping rules.
Figure 3: Cost-adjusted simple regret across acquisition–stopping rule pairs in 8D Bayesian regret experiments. Objectives are sampled from a GP with a Matérn-5/2 kernel, using three different cost functions scaled by $\lambda=0.01$. The PBGI/LogEIPC stopping rule consistently yields the lowest cost-adjusted regret that is close to hindsight optimal, when paired with LogEIPC, PBGI, and LCB.
4.2 Empirical Automated Machine Learning Benchmarks
@@
We then evaluate on two empirical benchmarks based on use cases from hyperparameter optimization and neural architecture search: LCBench (Zimmer et al., 2021) and NATS-Bench (Dong et al., 2021).
@@
LCBench provides training data of 2,000 hyperparameter configurations evaluated on multiple OpenML datasets. Due to the space constraints, we report results on the first three datasets from the lite version, covering image, mixed-type tabular, and large-scale numerical tabular classification tasks. In the unknown-cost experiments, the cost is the full (200-epoch) training time. For the known-cost setting, we observe that training time to scale approximately linearly with the number of model parameters. Based on a linear regression fit (see Appendix˜C), we adopt $10^{-3}\times p$, where $p$ is the number of model parameters, as a proxy of training time.
@@
NATS-Bench provides a search space of 32,768 neural architectures varying in channel sizes across layers, evaluated on three datasets. As in LCBench, for the known-cost experiments, we approximate the full runtime (training + validation time) at 90 epochs using $\alpha F+\beta$, where $F$ is the number of floating point operations; see Appendix˜C for details.
@@
We report the mean and error bars (2 times standard error) of the cost-adjusted simple regret, where we consider the evaluation costs to be some representative cost-scaling factor $\lambda$ (see Section˜3.2) multiplied with cumulative runtime. Simple regret is computed as the difference between (a) test error of the configuration with best validation error at the stopping time and (b) the best test error over all configurations. We present the results using proxy runtime here, and defer to Appendix˜D the additional results under (i) the cost model mismatch scenario (proxy runtime used during Bayesian optimization but actual runtime used for evaluation), and (ii) the unknown-cost scenario (actual runtime used during Bayesian optimization).
@@
Figure˜4 presents cost-adjusted regret for combinations of acquisition function and stopping rule on classification tasks from LCBench and NATS-Bench, with representative [^3] These were chosen to avoid degenerate cases—neither so large that the policy stops after only a few evaluations (e.g., $\lambda=10^{-4}$ on NATS-Bench, Figure 18) nor so small that evaluations are effectively free. cost-scaling factor $\lambda=10^{-4}$ and $\lambda=10^{-5}$ respectively. Additional results under different cost-scaling factors and the unknown-cost setting are provided in Figures˜15, 16 and 17 of Appendix˜D, along with visual comparisons between adaptive and fixed-iteration stopping. Overall, our PBGI/LogEIPC stopping rule consistently performs strongly in terms of cost-adjusted simple regret, particularly when paired with the PBGI acquisition function. We also report, in Table˜1 of Appendix˜D, how often each stopping rule fails to trigger within the 200-iteration cap: baselines such as SRGap-med and UCB-LCB often fail to stop early—particularly on NATS-Bench—whereas our stopping rule reliably stops before the cap.
@@
As we transition from model match to model mismatch, where the true objective function does not align perfectly with the Gaussian process prior, we find that our stopping rule, when paired with the PBGI acquisition function, continues to deliver performance close to the hindsight optimal, except on two NATS datasets likely affected by severe mismatch. In contrast, pairing with the LogEIPC acquisition function is less competitive. This degradation appears to stem from the relative advantage of PBGI over LogEIPC in higher dimensions or misspecified settings: as shown in Figure˜11 in Appendix˜D, PBGI maintains stronger performance under misspecification, thus contributing to the improved overall cost-adjusted regret. Thus, while our stopping rule is broadly robust, the choice of acquisition function remains an important consideration when facing objective model mismatch.
Figure 4: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on LCBench and NATS-Bench. The objective is to minimize validation error on classification tasks, with proxy runtime as evaluation cost, scaled by representative values of $\lambda$ ($10^{-4}$ for LCBench and $10^{-5}$ for NATS-Bench); see Figures˜12, 13 and 14 for performance under other $\lambda$. The matched PBGI combination performs best on adult, higgs, and cifar10-valid, second-best on Fashion-MNIST, and third-best on the remaining two, while closely approaching the hindsight optimal across all tasks. Our stopping rule is also competitive when paired with LogEIPC or TS.